OpenStudy (anonymous):

What is the Least Common Denominator (LCD) of eleven-fourteenths and sixteen twenty-firsts

OpenStudy (asnaseer):

you just need to find the lowest common multiple of 14 and 21. you can write these as the product of their primes as follows: 14 = 7 * 2 21 = 7 * 3 so least common multiple = 7 * 2 * 3